Consider the following reaction: 2 HCl + CaCO 3 → CaCl 2 + H 2 O + CO 2

How many mols of calcium chloride can be produced if you begin with 7.49 mL of 0.62 M HCl and 12.26 grams of calcium carbonate? Record your answer in scientific notation, using 3 significant figures.
Answer #1

Moles of calcium carbonate used = mass/molar mass

= 12.26/100.08 = 0.1225 moles

Moles of HCl = molarity×volume/1000

= 0.62×7.49/1000 = 0.00464 moles

According to the reaction,

2mole HCl require 1 mole CaCO3

0.00464 moles of HCl require 0.00464/2 moles CaCO3

= 0.00232 moles CaCO3

We have more than this CaCO3 hence HCl is limiting reagent.

According to the reaction,

2 mole HCl give 1 mole CaCl2

0.00464 moles HCl give 0.00464/2 mole CaCl2

= 0.00232 moles CaCl2